Your division is considering two investment projects, each of which requires an upfront expenditure of $15...

  1. Your division is considering two investment projects, each of which requires an upfront expenditure of $15 million. You estimate that the investments will produce the following cash flows:

                                                                                                                                    Year                   Project A                                                    Project B                                                       1                    $5,000,000                                             $20,000,000                                                    2                 $10,000,000                                             $10,000,000                                                       3                 $20,000,000                                             $6,000,000                                                                       

  1. What is each project’s Net Present Value, assuming that your division’s cost of capital is 5%? 10%? 15%?

       (b) What is each project’s Internal Rate of Return?
Answer #1

Calculate the NPV and IRR as follows:

А B 141 Year 142 143 144 c Project A Project B 0 -$15,000,000.00 -$15,000,000.00 $5,000,000.00 $20,000,000.00 2 $10,000,000.0

Formulas:

141 Year 142 0 143 1 144 2 145 3 146 NPV @ 5% 147 NPV @ 10% 148 NPV @ 15% 149 IRR В Project A Project B =-15*1000000 =-15*100